OpenStudy (anonymous):

Find the values of the six trigonometric functions for an angle A in standard position if a point with coordinates (2, 0) lies on its terminal side.

OpenStudy (anonymous):

Do you have a drawing or do I need to make one?

OpenStudy (anonymous):

i dont have one! @MrFrizzy

OpenStudy (mayankdevnani):

values of the six trigonometric functions of an angle in standard position if the point with coordinates (6, 8) lies on its terminal side. In rightt angle triangle height is 0 and base is 2, So hypotenuse is ??

OpenStudy (mayankdevnani):

can you find hypotenuse?

OpenStudy (anonymous):

2?

OpenStudy (anonymous):

what does (6,8) have to do with this? @mayankdevnani

OpenStudy (mayankdevnani):

ok.

OpenStudy (mayankdevnani):

6 and 8 are sides of right angle triangle

OpenStudy (mayankdevnani):

can i go on!!

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

but not this right angle triangle right?

OpenStudy (mayankdevnani):

sorry...not 6,8 its 2,0 sorry @kenneyfamily

OpenStudy (anonymous):

ok got it

OpenStudy (mayankdevnani):

In right angle triangle height is 0 and base is 2, So hypotenuse is 2 sin= 0/2 cos= 2/2 tan= 0/2 csc= 2/0 sec= 2/2 cot=2/0

OpenStudy (mayankdevnani):

ok @kenneyfamily

OpenStudy (anonymous):

THANKS @mayankdevnani

OpenStudy (mayankdevnani):

welcome